# Krylov Complexity in early universe

### v2

@@ Abstract
The Lanczos algorithm offers a method for constructing wave functions for both closed and open systems based on their Hamiltonians. Given that the entire early universe is fundamentally an open system, we apply the Lanczos algorithm to investigate Krylov complexity across different phases of the early universe, including inflation, the radiation-dominated period (RD), and the matter-dominated period (MD). Notably, we find that Krylov complexity differs between the closed and open system approaches. To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials, including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential, which is taking into account the violations of slow-roll conditions. This analysis is conducted in terms of conformal time through the preheating process. Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar within distinctive potentials in RD and MD. Additionally, we rigorously construct what is referred to as an open two-mode squeezed state, utilizing the second kind of Meixner polynomials. Based on this construction, we are the first to calculate the evolution equations for $r_{k}$ and $\phi_{k}$ as they relate to the scale factor. Our findings suggest that dissipative effects lead to a rapid decoherence-like behavior. Moreover, our results indicate that inflation behaves as a strongly dissipative system, while both the radiation-dominated and matter-dominated phases exhibit characteristics of weak dissipation. This research provides new insights into exploring the universe from the perspective of quantum information.
1 Introduction
The idea of complexity was initially introduced in the holographic framework Stanford:2014jda . In this framework, it was observed that the conformal field theory (CFT) reaches thermal equilibrium relatively quickly, while the corresponding bulk gravity (wormhole) takes a much longer time to reach the same state. To address this issue, Refs. Hartman:2013qma ; Liu:2013iza introduced the concept of computational complexity within the holographic framework. This concept suggests that while the emergence of spacetime is related to quantum entanglement, the evolution of space occurs through computational complexity.
The concept of complexity has become increasingly essential in high-energy physics even without the holographic principle. First, the so-called computational complexity describes the minimum number of logical steps required for a task Aaronson:2016vto . Furthermore, the method of calculating the complexity was first proposed by Nielson, $\it e.t.c$, dubbed as the geometrical method Nielsen:2005mkt ; Dowling:2006tnk ; Nielsen:2006cea . Another approach is highly related to the ”Fubini study” distance of the geometry of information Chapman:2017rqy . These two geometrical methods are highly dependent on the choice of the geometry of the parametric manifold. For more precisely capturing the information for the quantum state itself, the concept of Krylov complexity was proposed using the Lanczos algorithm viswanath1994recursio , which is explicitly related to the quantum state without needing the parametric manifold. Since this method is free of ambiguities compared with the previous two ones, there will be similarities and differences between them. According to the concept of Krylov complexity, the universal operator growth was proposed Parker:2018yvk . Once the precise quantum state is given, we could analyze its Krylov complexity, Krylov entropy, and Lanczos coefficient based on the Lanczos algorithm viswanath1994recursio . For the connection between the Krylov complexity and the circuit complexity, Ref. Aguilar-Gutierrez:2023nyk has shown that they cannot be compatible with each other. However, Ref. Caputa:2021sib demonstrated that the Krylov complexity was proportional to the distance of the ”Fubuni-Study” metric.
Recently, the Krylov complexity has been broadly applied to condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically analyzed the Krylov complexity in various polynomials dubbed as the orthogonal basis. Due to the importance of the SKY model in high energy physics, the Krylov complexity can also be implemented into this field Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk . Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp . Ref. Dymarsky:2019elm proved that the chaos can be understood as a delocalization in Krylov complexity. Furthermore, one could observe the exponential growth of Krylov complexity in integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt . More recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj ; Balasubramanian:2024ghv ; Bhattacharya:2024hto ; Mohan:2023btr . Especailly, the concept of Krylov complexity can be extended into in various fields Nandy:2024htc .
We aim to investigate the Krylov complexity for the whole early universe, including the inflationary period, radiation-dominated period (RD), and matter-domination period (MD). Some research has examined the computational complexity for inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked , where the analysis demonstrated that there is oscillation at the very beginning of inflation, and the evolution was similar in post inflation period. Once considering the quantum effects via various quantum frameworks, the complexity will show the irregular oscillation during inflation providing a new method for distinguishing various inflationary models Li:2023ekd . Quite a few researches are investigating the Krylov complexity in the early universe. Ref. Adhikari:2022oxr is the first one who investigate Krylov complexity in inflation by evaluating the values of non-trivial sound speed, where their method is based on the closed system’s method. However, our universe is an open system. In light of the open system’s method, Ref. Li:2024kfm investigated the Krylov complexity within the framework of modified dispersion relation in inflation, indicating that the inflation behaves as a strong dissipative system and the overall trend of Krylov complexity is increasing. The key ingredient for exploring the Krylov complexity with the open system’s method is constructing the wave function in light of Ref. Bhattacharya:2022gbz , whose wave function wave from Li:2024kfm will serve as our starting point in this work.
In previous studies, Krylov complexity was applied to the early universe using the thermal state. Refs Li:2024ljz ; Li:2024iji examined Krylov complexity by considering the contribution of the potential (quadratic term) and the violation of slow-roll conditions. The potential was approximated to be the quadratic term around the minimum in our previous work Li:2024ljz , which may not capture fully precise information. In the upcoming work, we plan to include more models such as the Higgs potential Higgs:1964pj , the potential of Starobinsky inflation ($R^{2}$ inflation) Starobinsky:1980te , and the potential of chaotic inflation Linde:1983gd . These models can cover most inflationary models, especially those favored by current observations, which indicate that the potential shape is a plateau. It has been suggested by Ref. Galante:2014ifa that many inflationary models can be unified into the so-called $\alpha$ attractors, whose potential is similar to $R^{2}$ inflationary potential. Additionally, Ref. Albrecht:1992kf demonstrated a relation between the Bunch-Davies vacuum and the two-mode squeezed state. We will continue to use the two-mode squeezed state to investigate the Krylov complexity for the entire early universe in light of these two factors.
@@ 1 Introduction
This paper will be organized as follows. In Sec. 2, we will introduce the Lanczos algorithm. Sec. 3 will give a brief introduction to the early universe, including inflation, RD, and MD, and three different kinds of potentials. Sec. 4 will give the two-mode squeezed state, and the numerical solutions of $r_{k}$ in different periods. Sec. 5 will investigate the Krylov complexity in a method of closed system. Sec. 5.2 will investigate the Krylov entropy in a method of closed system. Sec. 6 will give the Lanczos coefficient and Krylov complexity with the approach of open system, where we are the first to derive the evolution equation of $\phi_{k}$ and $r_{k}$ based on the open two-mode squeezed state. Finally, the conclusion and outlook will be given in Sec. 7.
2 Lanczos algorithm with the approach of closed system
@@
In this section, we will mainly introduce the Lanczos algorithm with the approach of closed system. First, the operator $\mathcal{O}(t)$ in Heisenberg picture needs to be defined as follows,
@@
where H is the Hamilton and its solution is
@@
Next, let’s define the Liouvillian super-operator $\mathcal{L}_{X}$ as $\mathcal{L}_{X}Y=[X,Y]$, where [,][,][ , ] denotes the commutator. We apply this to Eq. (2), the operator can also be expressed as
@@
$\begin{split}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}),\mathcal{L}^{1}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{1}),\mathcal{L}^{2}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{2}),\mathcal{L}^{3}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{3})...\end{split}$ (4)
However, there are not orthogonal bases. In order to construct an orthogonal basis, we use the Lanczos algorithm. The first two operators become
@@
where $b_{1}=\sqrt{(\tilde{\mathcal{O}}_{0}\mathcal{L}|\mathcal{L}\tilde{\mathcal{O}}_{0})}$ desribing the normalized vector. And the next orthogonal basis can be written by
@@
where $b_{n}$ represents the Lanczos coefficients. This iterative relation will stop and produce the finite orthogonal Krylov basis if $b_{n}=0$. Subsequently, we can express Eq. (3) as follows,
@@
Next, we can define the Krylov complexity using the wave function $\phi_{n}$ as follows
@@
Additionally, the Lanczos coefficient can be bounded via Parker:2018yvk ,
$b_{n}\leq\alpha n+\eta$ (11)
where $\alpha$ and $\eta$ encodes different information in various models. It should be noticed that $b_{n}=\alpha n+\eta$, meaning that the system is a maximal chaotic dynamical system. For the case of maximal chaos, we could build the relation between the Lyapunov index and $\gamma$ and $\alpha$ as follows,
@@
3 Some setup of early universe
In this section, we will first give some basis of scale factor in distinctive periods. Following the notation of our previous work Li:2024ljz , one of the core goal of this work is examining the impact of inflationary potential for the Krylov complexity in RD and MD due to the violation of slow-roll conditions. The Mukhannov-Sasaki variable contains the contribution of potential which cannot show the explicit contribution to the Krylov complexity. That is why we just perturb the inflaton field explicitly in this paper. Thereafter, we will give some very important inflationary potentials, namely, the chaotic inflationary potential, Starobinksy’s inflationary potential ($R^{2}$ inflation) and Higgs potential.
3.1 Inflation, RD and MD
We will follow the notation of Baumann:2009ds to introduce the scale factor. The background metric is the Friedman-Lemaitre-Robertson-Walker metric,
@@ 3.1 Inflation, RD and MD
where $a(\eta)$ is the scale factor. Here, we will assume a single component universe, which means that the universe will experice the inflation, RD and MD, respectively. We could use equation of state to characterize the universe via
@@
where $P_{I}$ and $\rho_{I}$ is the pressure and energy density of the corresponding periods parameterized by $I$. For obtaining the relation between the conformal time and the scale factor, we need the comoving distance,
@@
where $\chi_{ph}(\eta)$ is the comoving distance in the event horizon of particle. In light of $w_{I}$, the Hubble radius can be replaced by
@@
where we use $\omega$ to take place of $\omega_{I}$ for simplify. Integrating equation (15) yields the relation between the scale factor and the conformal time
$\eta=\frac{2H_{0}^{-1}}{(1+3\omega)}a^{\frac{1}{2}(1+3\omega)}$ (17)
Consequently, we can derive the relation for various time periods
@@
One could see that the various values of $\omega$ describe the distinctive periods.
In the following text, we will discuss the fundamental basis for the different periods in the evolution of the universe. Initially, the universe is created by a singularity, followed by a period of exponential expansion known as inflation. During this phase, the curvature perturbation transitions from the quantum level to the classical level after the horizon exits, leaving an imprint on the Cosmic Microwave Background (CMB). Subsequently, the universe enters the Radiation Dominated (RD) era, characterized by an energy density given by the equation $\rho_{\rm rad}(t)=\rho_{\rm rad_{0}}\frac{a_{0}^{4}}{a(t)^{4}}$, where $\rho_{\rm rad_{0}}$ and $a_{0}$ represent the initial values of the energy density and the scale factor of RD. During the RD era, particles are relativistic, but as the temperature decreases, they become increasingly non-relativistic. This leads the universe into the matter-dominated (MD) epoch, where the energy density is given by $\rho_{\rm mat_{0}}(t)=\rho_{\rm mat_{0}}\frac{a_{0}^{3}}{a(t)^{3}}$. The transition from the RD to the MD epoch occurs when the energy density of radiation becomes equal to the energy density of matter. Once entering MD, there is a process called preheating where the particles will be generated via this process. First, we give the full action for the single field inflation,
[…]
@@
where $\ddot{\phi}(t)=\partial_{t}^{2}\phi(t)$ and the similar work can be found in our previous works Liu:2019xhn ; Liu:2021rgq . Defining with the effective mass $m_{\rm eff}=\frac{d^{2}V}{d\phi^{2}}$ and changing variable as $\tilde{\phi}(t)=a^{3/2}\phi(t)$, Eq. (20) will become
@@
the solution of Eq. (21) is $\phi(t)=a^{-3/2}\cos(m_{\rm eff}t)$ where it is the solution of $\phi(t)$ in preheating period and we have neglecte the term of $\frac{9}{4}H^{2}\tilde{\phi}+\frac{3}{2}\dot{H}\tilde{\phi}$ since it is much smaller compared with potential. It should be noticed that the solution is in terms of physical time $t$. In this paper, we will use $d\eta=\frac{dt}{a}$ to transform into the conformal time for the latter investigations, which is denoted by Eq. (18).
3.2 Various inflationary potentials
There are lots of inflationary models that represent distinctive inflationary potentials. Current observations favor the concave potentials Planck:2018jri , where the potential of $R^{2}$ inflation belongs to this potential. Additionally, the $R^{2}$ inflation agrees with the observations from inflation to large scales. From the theoretical aspect, we could see that many inflationary models can be unified into the so-called $\alpha$ attractors inflationary models Galante:2014ifa , where the other models and $R^{2}$ inflationary model belong to the same kind of inflationary models. Being armed with these two reasons, we will introduce $R^{2}$ inflation. According to the current standard model, the Higgs field is the only scalar field in nature. Therefore, the Higgs potential is our other option Higgs:1964pj . The final choice is the chaotic inflation whose potential is the quadratic term Linde:1983gd . Finally, we could see that our choices could cover the most inflationary models.
Here, we list the formulas of distinctive inflationary potentials as follows,
$\begin{cases}&\ V_{1}=\frac{1}{2}m^{2}\phi^{2},\\ &\ V_{2}=\frac{3}{4}M^{2}M_{p}^{2}(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}})^{2},\\ &\ V_{3}=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4},\end{cases}$ (22)

### v3

@@ Abstract
The Lanczos algorithm offers a framework for constructing wave functions in closed and open quantum systems from their Hamiltonians. Since the early universe is inherently an open system, we employ this algorithm to investigate Krylov complexity across various cosmological phases: inflation, radiation domination (RD), and matter domination (MD). Our results highlight a clear distinction in Krylov complexity between the closed- and open-system methodologies. To accurately capture the influence of potentials during RD and MD, we examine a set of inflationary potentials, including the Higgs potential, $R^{2}$ inflation, and chaotic inflation, while incorporating violations of slow-roll conditions. This study is conducted in conformal time through the preheating stage. Numerically, we find that the evolution of Krylov complexity and Krylov entropy shows remarkable similarity across different potentials during RD and MD. Furthermore, we rigorously construct an open two-mode squeezed state using the second kind of Meixner polynomial. Based on this construction, we derive for the first time the evolution equations for the squeezing parameter $r_{k}$ and phase $\phi_{k}$ in terms of the scale factor. Our analysis indicates that dissipative effects lead to rapid decoherence-like behavior. In addition, we observe that the inflationary universe behaves as a strongly dissipative system, whereas during the RD and MD epochs the universe exhibits weak dissipative characteristics. This work opens new perspectives for studying the universe from a quantum-informational viewpoint.
I Introduction
The concept of complexity has become increasingly important in high-energy physics, although a unified definition is still lacking. Nonetheless, several computational frameworks have been proposed to evaluate it. One major approach is computational complexity, which quantifies the minimum number of logical steps required to perform a given task Aaronson:2016vto . Another significant line of research stems from Nielsen and collaborators, who introduced a geometric method for assessing complexity Nielsen:2005mkt ; Dowling:2006tnk ; Nielsen:2006cea . A related framework connects complexity to the Fubini–Study distance in the geometry of information space Chapman:2017rqy . However, both geometric approaches depend strongly on the choice of metric on the underlying parameter manifold.
A more recent development is the notion of Krylov complexity, which circumvents the need to construct a parameter manifold altogether viswanath1994recursio . By operating directly in the operator Hilbert space, this approach avoids certain ambiguities inherent in geometric formulations, though the precise connections and distinctions between Krylov complexity and other frameworks remain an active area of research. Within this paradigm, the concept of universal operator growth has been proposed Parker:2018yvk , suggesting that the late-time dynamics of quantum operators exhibit universal features encoded in their Krylov complexity. Given a specific quantum operator, one can explicitly compute its Krylov complexity, associated Krylov entropy, and the sequence of Lanczos coefficients using the recursive Lanczos algorithm viswanath1994recursio . The relationship between Krylov complexity and circuit complexity has been examined from contrasting viewpoints: while Ref. Aguilar-Gutierrez:2023nyk argues that the two notions are generally incompatible. However, Ref. Caputa:2021sib shows that, under certain conditions, Krylov complexity is proportional to the distance measured by the Fubini–Study metric.
Recently, Krylov complexity has found broad applications across both condensed matter and high-energy physics. A systematic study employing various orthogonal polynomial bases was carried out in Ref. Muck:2022xfc , highlighting the flexibility and robustness of the Krylov framework. Given the central role of the Sachdev–Ye–Kitaev (SYK) model in understanding quantum chaos and holography, Krylov complexity has been actively explored in this context as well Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk . Beyond the SYK model, it has been applied to the analysis of generalized coherent states Patramanis:2021lkx , spin systems such as the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , and conformal field theories Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk . More recently, the formalism has been extended to topological phases of matter Caputa:2022eye and even to quantum field theory He:2024hkw ; He:2024xjp , underscoring its growing relevance as a universal diagnostic of operator dynamics and quantum complexity.
It was shown in Ref. Dymarsky:2019elm that quantum chaos can be interpreted as delocalization in Krylov space, an insight that has reshaped the understanding of operator growth in chaotic systems. Interestingly, exponential growth of Krylov complexity has also been observed in certain integrable models exhibiting saddle-dominated scrambling, with universal features confirmed in Ref. Huh:2023jxt . Since then, the scope of Krylov complexity has continued to expand rapidly. Recent developments include its application to open quantum systems, non-Hermitian dynamics, holography, random matrix theory, and out-of-equilibrium phenomena—see, e.g., Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj ; Balasubramanian:2024ghv ; Bhattacharya:2024hto ; Mohan:2023btr . Notably, the framework has also been generalized beyond standard unitary quantum mechanics, with extensions to stochastic dynamics, classical systems, and information-theoretic settings Nandy:2024htc .
In this work, we aim to investigate Krylov complexity throughout the entire evolution of the early universe, spanning the inflationary epoch, the subsequent RD era, and the MD era. Previous studies have explored computational complexity during inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked , uncovering an oscillatory behavior at the onset of inflation and a qualitatively similar pattern in the post-inflationary phase. When quantum effects are incorporated through various quantum-information-theoretic frameworks, the complexity displays irregular oscillations during inflation, a feature that provides a novel diagnostic for distinguishing between competing inflationary models Li:2023ekd .
While several recent works have begun to explore Krylov complexity in the context of the early universe, Ref. Adhikari:2022oxr stands out as the first to investigate it during inflation, specifically accounting for non-trivial sound speed effects within a closed quantum system framework. However, given that the early universe is more realistically modeled as an open quantum system, interacting with unobservable environmental degrees of freedom, a more physically complete treatment requires going beyond unitary evolution. Addressing this, Ref. Li:2024kfm adopted an open quantum system approach to study Krylov complexity under a modified dispersion relation during inflation. Their analysis reveals that inflation behaves as a strongly dissipative process, with Krylov complexity displaying a persistent overall increase. A crucial ingredient in this open-system formulation is the explicit construction of the time-evolving wave function, as systematically developed in Ref. Bhattacharya:2022gbz . In the present work, we take the wave function derived in Ref. Li:2024kfm as our starting point to extend the analysis across the full cosmological timeline, from inflation through RD and MD eras.
In previous studies, Krylov complexity has been applied to the early universe using a thermal state as the reference. Refs. Li:2024ljz ; Li:2024iji extended this framework by incorporating the contribution of the scalar potential—approximated by its quadratic term, and examined the impact of slow-roll violation on complexity growth. In our earlier work Li:2024ljz , the potential was expanded quadratically around its minimum; while analytically convenient, this approximation may not fully capture the dynamics of realistic inflationary scenarios.
It has been suggested by Ref. Galante:2014ifa that many inflationary models are unified into the framework of $\alpha$-attractors, whose potential form closely resembles that of $R^{2}$ inflation. Moreover, Ref. Albrecht:1992kf established a connection between the Bunch-Davies vacuum and the two-mode squeezed state. Motivated by these two key insights, we will continue to employ the two-mode squeezed state formalism to study Krylov complexity in the context of the entire early universe. To improve physical fidelity, we plan in upcoming work to generalize the analysis to well-motivated inflationary potentials, including the Higgs potential Higgs:1964pj , Starobinsky’s inflation Starobinsky:1980te , and chaotic inflation Linde:1983gd . These models encompass most inflationary models and are especially relevant in light of current observational constraints, which strongly favor plateau-like potentials. This extension will allow for a more accurate and observationally grounded exploration of quantum complexity during inflation.
@@ I Introduction
This paper is organized as follows. In Sec. II, we will introduce the Lanczos algorithm. Sec. III will provide a brief review of the early universe, covering inflation, RD, and MD eras, along with three representative inflationary potentials. In Sec. IV, we will present the two-mode squeezed state formalism and numerically solve for the parameter $r_{k}$ across different cosmological periods. Sec. V will investigate the Krylov complexity in the framework of a closed-system methodology, which Sec. V.2 will examine the corresponding Krylov entropy. In Sec. VI, we will analyze the Lanczos coefficients and Krylov complexity using an open-system methodology; notably, we will present the evolution equations for $\phi_{k}$ and $r_{k}$ derived within the open two-mode squeezed state formalism in the first time. Finally, conclusions and outlook will be given in Sec. VII.
In this work, we will work in Planck units, namely $c=M_{P}=G=1$.
II Lanczos algorithm with closed-system methodology
@@
In this section, we will mainly review the Lanczos algorithm with closed-system methodology. First, the operator $\mathcal{O}(t)$ in Heisenberg picture needs to be defined as follows,
@@
where H is the Hamiltonian and its solution is
@@
Next, let’s define the Liouvillian super-operator $\mathcal{L}_{X}$ as $\mathcal{L}_{X}Y=[X,Y]$, where $[,]$ denotes the commutator. We apply this to Eq. (2), the operator is also expressed as
@@
[…]
However, they are not orthogonal bases. In order to construct an orthogonal basis, we use the Lanczos algorithm. The first two operators become
@@
where $b_{1}=\sqrt{(\tilde{\mathcal{O}}_{0}\mathcal{L}|\mathcal{L}\tilde{\mathcal{O}}_{0})}$ describing the normalized vector. And the next orthogonal basis is written by
@@
where $b_{n}$ represents the Lanczos coefficients. This iterative relation will stop and produce the finite orthogonal Krylov basis if $b_{n}=0$. Subsequently, we express Eq. (3) as follows,
@@
Next, we define the Krylov complexity using the wave function $\phi_{n}$ as follows
@@
Additionally, the Lanczos coefficient is bounded via Parker:2018yvk ,
$b_{n}\leq\alpha n+\eta,$ (11)
where $\alpha$ and $\eta$ encode different information in various models. Note that $b_{n}=\alpha n+\eta$, which indicates the system is a maximally chaotic dynamical system. For the case of maximal chaos, we obtain the relationship between the Lyapunov exponent and $\gamma$ and $\alpha$ as follows
@@
III Some setup of early universe
In this section, we first review some fundamental aspects of the scale factor in different cosmological epochs. Following the notation of our previous work Li:2024ljz , a central goal of this study is to examine the impact of the inflationary potential on Krylov complexity during the RD and MD eras, particularly due to the violation of slow-roll conditions. The Mukhanov-Sasaki variable incorporates contributions from the potential in a way that does not explicitly reveal its influence on Krylov complexity. Therefore, in this work, we perturb the inflaton field directly. We then introduce several important inflationary potentials: the chaotic inflationary potential, the Starobinsky potential ($R^{2}$ inflation), and the Higgs potential. Finally, we employ preheating dynamics to numerically evaluate the effective mass of the inflaton with respect to the scale factor.
III.1 Inflation, RD and MD
We will follow the notation of Baumann:2009ds to introduce the scale factor. The background metric is the Friedman-Lemaitre-Robertson-Walker (FLRW) metric,
@@ III.1 Inflation, RD and MD
where $a(\eta)$ is the scale factor expressed in conformal time $\eta$. Alternatively, the FLRW metric can also be written in terms of physical time $t$ as $ds^{2}=-dt^{2}+a(t)^{2}d\vec{x}^{2}$, from which the relation $dt=ad\eta$ follows explicitly. Here, we consider a universe dominated by a single component during each epoch, meaning it undergoes successive phases of inflation, RD, and MD. The evolution of the scale factor in each era can be characterized by the equation of state parameter $w$ via
@@
where $P_{I}$ and $\rho_{I}$ are the pressure and energy density, respectively, in each epoch $I$. To derive the relation between conformal time and the scale factor, we make use of the comoving distance,
@@
where $\chi_{\mathrm{ph}}(\eta)$ is the comoving particle horizon distance, and we have used the definition of the Hubble parameter $H=\dot{a}/a$, with $\dot{a}=da/dt$. In terms of the equation of state parameter $w_{I}$, the Hubble radius can be expressed as
@@
where, for simplicity, we replace $\omega_{I}$ with $\omega$. Integrating Eq. (15) yields an explicit functional relation between the scale factor and the conformal time,
$\eta=\frac{2H_{0}^{-1}}{(1+3\omega)}a^{\frac{1}{2}(1+3\omega)}.$ (17)
Consequently, we derive the relation for various time periods
@@
It shows that the various values of $\omega$ describe the distinctive periods.
In the following, we outline the fundamental evolution of the universe across its distinct epochs. The universe emerges from a singularity and undergoes a phase of exponential expansion known as inflation. During this period, curvature perturbations transition from quantum to classical after horizon exit, becoming imprinted on the Cosmic Microwave Background (CMB). Subsequently, the universe enters the RD era, characterized by an energy density $\rho_{\rm rad}(t)=\rho_{\rm rad_{0}},a_{0}^{4}/a(t)^{4}$, where $\rho_{\rm rad_{0}}$ and $a_{0}$ denote the initial energy density and scale factor at the onset of RD. In this epoch, particles are relativistic. As the temperature drops, particles become increasingly non-relativistic, leading the universe into the MD era, where the energy density evolves as $\rho_{\rm mat}(t)=\rho_{\rm mat_{0}},a_{0}^{3}/a(t)^{3}$. The transition from RD to MD occurs when the energy densities of radiation and matter equalize. During the subsequent MD epoch, a preheating process can generate particles. We begin by presenting the full action for single-field inflation:
$S=\int d\eta d^{3}\vec{x}\sqrt{-g}\bigg[\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V(\phi)\bigg],$ (19)
@@
where $\ddot{\phi}(t)=\partial_{t}^{2}\phi(t)$, $V(\phi),_{\phi}=\frac{dV(\phi)}{d\phi}$ and the similar work is found in our previous works Liu:2019xhn ; Liu:2021rgq . Defining with the effective mass as follows,
$m_{\rm eff}=\frac{d^{2}V}{d\phi^{2}}=V_{,\phi\phi}.$ (21)
Here, we explain why the effective mass is defined as $m_{\rm eff}=d^{2}V/d\phi^{2}$. For the different inflationary potentials in Eq. (23), we take $V_{\rm chaotic}$ as an example, where one can directly show that $m_{\rm eff}^{2}=m^{2}$. Based on this, we adopt the same definition of effective mass for all inflationary potentials considered. To solve Eq. (20), we introduce a new variable $\tilde{\phi}(t)=a^{3/2}\phi(t)$, which transforms Eq. (20) into
@@
The solution to Eq. (22) is $\phi(t)\propto a^{-3/2}\cos(m_{\rm eff}t)$, which describes the behavior of $\phi(t)$ during the preheating period. Here, the terms $\frac{9}{4}H^{2}\tilde{\phi}+\frac{3}{2}\dot{H}\tilde{\phi}$ have been neglected, as they are significantly smaller than the potential contribution. Note that this solution is expressed in physical time $t$. For the subsequent analysis in this paper, we will transform it into conformal time using $d\eta=dt/a$, as given by Eq. (18).
III.2 Various inflationary potentials
In this subsection, we will use Eq. (22) to establish a relation between the scale factor and various inflationary potentials. There exist many inflationary models, each characterized by a distinct potential. Current observational data favor concave potentials Planck:2018jri , a category that includes the potential of $R^{2}$ inflation. Furthermore, $R^{2}$ inflation remains consistent with observations from inflationary to large scales. From a theoretical perspective, many inflationary models can be unified within the framework of $\alpha$-attractor models Galante:2014ifa , which also encompass $R^{2}$ inflation. Motivated by these two considerations, we include $R^{2}$ inflation in our analysis. According to the Standard Model, the Higgs field is the only fundamental scalar field in nature; therefore, the Higgs potential serves as another natural choice Higgs:1964pj . Lastly, we consider chaotic inflation with a quadratic potential Linde:1983gd . Collectively, these choices cover a representative range of inflationary models.
We consider the following inflationary potentials:
$\begin{cases}&\ V_{\rm chaotic}=\frac{1}{2}m^{2}\phi^{2},\\ &\ V_{\rm R^{2}}=\frac{3}{4}M^{2}M_{p}^{2}(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}})^{2},\\ &\ V_{\rm Higss}=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4},\end{cases}$ (23)